# MiM: Mask in Mask Self-Supervised Pre-Training for 3D Medical Image Analysis

### v1

@@ 4.3.1 Comparison on the BTCV dataset
The scratch Swin-UNETR [28] only achieves 80.53% DSC. By pre-training with MiM on a 1k unlabeled dataset, we gained a 3.93% improvement with 84.46% DSC, which outperforms existing methods by a clear margin. Among the compared methods, Swin-UNETR [62] and MAE3D [16] achieved the best 82.58% and second-best 82.40% DSC, respectively. Our proposed MiM surpasses these two methods by 1.88% and 2.06% DSC, respectively, which is a clear improvement on this dataset.
It’s worth noting that scaling laws [41, 59, 20] also apply to 3D medical image pre-training. By pre-training with our larger scale unlabeled 10k dataset, we observed that Swin-UNETR and MAE3D achieved DSC scores of 83.20% and 83.52%, respectively. Our proposed MiM with 10k achieved a DSC score of 85.41%, which consistently outperformed these two methods by a large margin. These results suggest that scaling plays an important role in pre-training for 3D medical images and that our proposed MiM method is effective for pre-training on larger datasets.
Qualitative results. MiM was found to improve the completeness of segmentation results, as shown in Fig 5. The results of segmentation using MiM were better than those using Swin-UNETR [62], MAE3D [16], Unimiss [78], PCRLv2 [89] and GVSL [34] .
@@ 4.3.2 Comparison on the Unseen datasets
Promosing generalizability. We further conduct experiments on unseen datasets in pre-training, i.e., M-WHS [95], Spleen [58], Amos22 [37], and Flare22 [50]. The results of these four datasets are shown in Table 5. It can be observed that our MiM consistently outperformed all existing methods with a clear margin, which demonstrates promising generalizability to unseen datasets. Specifically, MiM outperformed existing methods by at least 1.89% DSC on average. By pre-training with a larger scale of unlabeled dataset 10k, Swin-UNETR [62] and MAE3D [16] improved by 0.92% and 0.94% DSC with 86.32% and 86.64%, respectively. Our MiM also gained improvement to 88.34% DSC and outperformed these two methods consistently.
Table 6: Experiment results on Task03 Liver, Task06 Lung, Task07 Pancreas, Task08 Hepatic Vessel and Task10 Colon on MSD dataset [58].
Method Pre-traning dataset Network Task03 Liver Task06 Lung Task07 Pancreas Task08 Hepatic Vessel Task10 Colon Avg
@@
Fig. 6: Visualization of segmentation results for four tasks on the MSD dataset [58]. We compared MiM with MAE3D [16], and Swin-UNETR [62]. The Dice Score (%) is provided for each segmented volume.
Qualitative results. To gain a better understanding of the improvements achieved by MiM, we present the visualization of segmentation results for four tasks on the MSD dataset [58]. We compared our MiM method with MAE3D [16] and Swin-UNETR [62]. It can be observed in Fig. 6 that MiM outperforms MAE3D [16] and Swin-UNETR [62], by effectively addressing the issues of under and over-segmentation outliers observed in the segmentation results obtained by these methods. Specifically, Swin-UNETR [62] exhibits under-segmentation due to the absence of certain parts of the labels such as the liver and hepatic vessel, leading to a lower DSC. On the other hand, in the case of the pancreas, over-segmentation is observed with MAE3D [16] and Swin-UNETR [62], where surrounding tissues are included with the label. In contrast, our proposed MiM method effectively delineates the boundary of the pancreas.
4.4 Comparison on the Covid-19-20 dataset
Generalization capacity on Lesion Segmentation task. To evaluate the generalizability of our MiM on the lesion segmentation task, we conducted experiments on the Covid-19-20 [82] dataset. Lesion segmentation involves identifying specific abnormalities, unlike organ segmentation that focuses on delineating anatomical structures [58, 82]. showed that our proposed MiM method consistently exhibited performance improvements compared to other SSL methods by at least 1.52% and 1.73% in terms of DSC and NSD, respectively, which is a clear improvement on this dataset. By pre-training with a larger scale of unlabeled dataset 10k, MiM further improved to 71.11% and 72.42% in DSC and NSD, respectively. These results demonstrate that MiM is effective for lesion segmentation tasks.
[…]
Method Pre-traning dataset Network Lesion segmentation
DSC(%) NSD(%)
From scratch
SegResNet [51] - - 61.38 61.73
3D UNet [55] - - 60.35 61.27
UNETR [29] - - 58.80 57.22
Swin-UNETR [28] - - 61.89 61.92
with General SSL
MoCov3 [15] 1k UNETR 63.15 64.32
localMIM [67] 1k UNETR 66.62 67.11
simMIM [79] 1k Swin-UNETR 67.22 68.19
[…]
with Medical SSL
Adam [35] 1k SegResNet 67.31 67.56
ModelGen [91] 1k 3D UNet 68.25 69.63
PCRLv1 [90] 1k 3D UNet 67.25 68.14
PCRLv2 [89] 1k 3D UNet 68.18 69.18
GL-MAE [94] 1k UNETR 67.11 68.94
[…]
Rubik++ [92] 1k Swin-UNETR 64.60 65.00
[…]
MiM 1k Swin-UNETR 70.14 71.02
[…]
4.5 Comparison on CC-CCII dataset
Generalization capacity on Disease Classification task. To evaluate the generalizability of our MiM on classification task, we fine-tuned it on the CC-CCII [82] dataset and compared its performance with state-of-the-art general and medical SSL methods. Since existing SSL methods didn’t conduct experiments on this dataset, we reproduced the related methods and reported the results. As shown in Table 8, it can be observed that MiM achieved the best performance in terms of ACC and AUC, surpassing all other methods with 93.63% and 99.39%. These findings demonstrate that the learned representation by MiM can be well transferable to classification problems and can be used effectively for medical image classification tasks. With a larger scale of pre-training dataset 10k, our MiM further improved to 94.12% and 99.52% in ACC and AUC, respectively, which shows the scalability of our proposed method when transferring across tasks.
@@ 4.5 Comparison on CC-CCII dataset
4.6 Comparison on the BraTS 21 dataset
Generalization capacity on MRI modality dataset. To evaluate the generalizability of our MiM on MRI datasets, we fine-tuned it on the BraST 21 [58] MRI tumor segmentation dataset and compared its performance with state-of-the-art general and medical SSL methods. It can be observed in Table 9 that both SSL methods can improve the performance of the model in segmenting tumors in BrasTS 21 [58] dataset. This is because CT and MRI are often used for the same task but for different purposes, and thus share similar anatomical structures. Therefore, the knowledge learned by SSL methods from unlabeled CT datasets can be transferred to MRI datasets [49, 62]. Our MiM outperformed all other methods by at least 1.34% with 79.28% DSC. By pre-training with a larger unlabeled dataset 10k, our MiM further improved to 79.92%, which shows the scalability of our proposed method when transferring across modalities.
[…]
@@ 4.6 Comparison on the BraTS 21 dataset
Label efficiency. The scarcity of labeled 3D medical images can pose a significant challenge, requiring models to perform well despite this constraint [62]. To evaluate the effectiveness of MiM under this constraint, we compared it with existing methods by reducing the number of manual labels at label ratios of 25% and 50%. As shown in Table 5. MiM outperformed all other methods by a significant margin when generalized to both BTCV [43] and unseen datasets. Specifically, MiM with 50% label achieved comparable performance with the scratch Swin-UNETR with 100% label with a clear margin.
4.6.1 Comparison on the MSD datasets
Generalizability to Organ Segmentation. To evaluate the generalizability, we conducted experiments on the five CT-based tasks on the MSD dataset [58], i.e., Task 03 Liver, Task06 Lung, Task07 Pancreas, Task08 Hepatic Vessel and Task10 Colon. Since existing methods didn’t conduct experiments with the same pre-training dataset, we reimplemented these methods for fair comparison. It can be observed in Table 6 that MiM achieves the best average DSC (70.07%) and NSD(78.75%) in all tasks. Since the scrtach Swin-UNETR [28] performs better UNETR [29] in terms of avg DSC (64.84% vs 62.98%) and NSD (73.12% vs 69.08). We further pre-trained UNETR [29] based on MiM for a fair comparison. It can be observed that with MiM pre-training, Swin-UNETR [28] gained an average improvement 5.23% and 5.63% in terms of DSC and NSD, respectively. With UNETR [29] as the network, we observed an average improvement of 10.09% and 9.29% in terms of DSC and NSD, respectively. Moreover, by pre-training with a larger scale of unlabeled dataset 10k, MiM further improved to 70.76% and 79.67% in DSC and NSD, respectively. These results demonstrate that MiM is effective for organ segmentation tasks.
@@ 4.7 Analysis of our proposed method
We first conducted ablation studies on loss functions and settings of MiM. Then, we conducted a hyper-parameter analysis for the loss terms. Lastly, we provided reconstruction results. All the models were pre-trained on 1k, which were then evaluated on BTCV [43] and MM-WHS [95].
@@ 4.7.1 Ablation study
Loss functions. To investigate the impact of hierarchical design by applying multi-level reconstruction and cross-level alignment, we conducted a comprehensive ablation study of the loss terms. Table 10 presents the results of the ablation studies on three validation datasets. By utilizing multi-level reconstruction, the model significantly improved upon the previous MiM method, as the model was able to capture multi-granularity visual cues of the 3D medical images and learn representation hierarchically, thereby learning broader information. The cross-level alignment for coarse $x^{1}$ and $x^{2}$ (4th row) achieved better performance than that of fine-cropped $x^{3}$ 3D medical images (5th row). Combining all losses could result in best performance improvements.
Table 10: Evaluation of the loss terms $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$. We report the performance on BTCV [43], MM-WHS [95], and Spleen [58] dataset.
@@
$\mathcal{L}^{\mathcal{R}}_{1}$ $\mathcal{L}^{\mathcal{R}}_{2}$ $\mathcal{L}^{\mathcal{R}}_{3}$ $\mathcal{L}^{\mathcal{C}}_{12}$ $\mathcal{L}^{\mathcal{C}}_{23}$ DSC(%) NSD(%) DSC NSD(%)
@@
Numbers of levels. We limit the number of levels $L$ to three. As illustrated in Fig. 7, we varied the number of levels $L$ and observed that the optimal performance was achieved when $L=3$. This choice stems from the fact that the resolution of the volume at Level-4 is considerably smaller in scale compared to the original volume ($1\times 1\times 1$), facilitating easier reconstruction but resulting in suboptimal performance [30, 79]. Moreover, the proximity of positive patches to negative patches within a confined region may hinder effective discrimination, particularly concerning hard negative patches [40]. Thus, we set the number of levels $L=3$ as the default choice.
Fig. 7: Evaluation of the number of level $L$ for our multi-level volume generation. We report the results on BTCV [43] and MM-WHS [95].
Reconstruction targets for $\mathcal{L}_{\mathcal{R}}$. In Fig. 3, we employ the mask patches from the coarser as the finer level volume, which is the default choice in MiM. In Table 11, we observed that adopting mask patches from the coarser level as the finer level volume results in better performance than the other choices.
Table 11: Evaluation of the process of finer level volume. We report the DSC on BTCV [43] and MM-WHS [95].
Finer level volume BTCV MM-WHS
@@
Visible patches 82.18 76.22 89.33 73.65
Maked patches 84.66 82.11 91.04 76.65
Negative pairs for $\mathcal{L}_{\mathcal{C}}$. In Eq. 6, we employ infoNCE loss[10] as the default choice. This loss maximizes the similarity between the cross-level images and pushes away the negative samples. Another loss function is BYOL cosine loss [24]. The primary distinction is whether to use negative samples. As demonstrated in Table 12, negative samples aid in learning better representations [12], and thus, we include them in our default choices.
Table 12: Evaluation of $\mathcal{L}_{C}$ whether using negative samples. We report the DSC on BTCV [43] and MM-WHS [95].
@@
Efficiency of MiM. Table 13 presents a comparative analysis of computation costs i.e. flops and times during the pre-training period across various methodologies. Our evaluation highlights the computational efficiency of MAE-based approaches when coupled with UNETR, as they leverage only the visible tokens from masked 3D medical images. In contrast, alternative MAE-based methods naively cooprating with hybrid backbones such as Swin-UNETR, necessitate processing all tokens. Our proposed method extends the hybrid backbone architecture [22] to 3D medical images in the pre-training stage, achieving significantly superior performance while maintaining computational efficiency.
Table 13: Evaluation of model complexity during pre-training and subsequent performance on downstream tasks. We report flops and pre-training duration for pre-training. We report Dice Score(%) on BTCV [43] and MM-WHS [95] dataset for fine-tuning.
Method Networks Tokens flops(G) Pre-training duration BTCV MM-WHS
All Partial
@@ 4.7.2 Hyper-parameter analysis
Weights of $\mathcal{L}_{\mathcal{R}}$ and $\mathcal{L}_{\mathcal{C}}$. In Eq. 10, the loss function of MiM consists of two parts, namely multi-level reconstruction and cross-level alignment. To determine the suitable class weight for these loss terms, we varied the value of $\alpha$ in Fig. 8. When we varied the value of $\alpha$ from a small value of $1e^{-2}$ to larger value $1e^{-1}$ during the pre-training period, it strengthened the impact of cross-level alignment between finer views and larger views, thereby enforcing the representation of local and global alignment. As a result, the DSC of the MiM fine-tuned on the BTCV validation dataset improved by $0.42\%$. However, further scaling up the weight of cross-level alignment did not bring any additional benefits. This may be because the magnitude of the cross-level alignment function is much larger than the magnitude of the reconstruction loss. Thus, scaling up the weight of cross-level alignment leads the model to ignore the reconstruction process, which is also crucial for MiM method.
Fig. 8: Evaluation of the loss weight $\alpha$ for $\mathcal{L_{C}}$. We report the results on BTCV dataset [43].
Weight of different levels of $\mathcal{L}_{\mathcal{R}}$. To determine the appropriate weights of each level for MiM, we applied different weights to determine the priority of the level of masked 3D medical images during the pre-training, as shown in Fig. 1. Specifically, we explored focusing on reconstructing the coarse level of masked images first, then the finer level (2nd row), and reconstructing the finer level first, then the coarse level of masked images (3rd row) by dynamically updating the weight with a cosine scheduler, as presented in Table 14. The results demonstrated that learning from finer to coarser masked images would help learn a more discriminative representation. Furthermore, we explored simultaneously learning the multi-granularity visual clues (4th row), which outperformed the other settings and set as default choice.
Table 14: Evaluation of the model trained with different learning paradiagm by varying the weight of each level of $\mathcal{L}_{\mathcal{R}}$, repectively. We reprot Disce Score(%) on BTCV [43] and MM-WHS [95] dataset.
Learning process Weight Datasets
Level-1 Level-2 Level-3 BTCV MM-WHS

### v2

@@ 4.2 Implementation Details
4.3 Experiments on downstream tasks
4.3.1 Comparison on the BTCV dataset
We first conducted experiments on the BTCV [53], and the results are presented in Table 3. Among the comparison methods, SimCLR [29], MoCov3 [66], DINO [49], localMIM [46], HPM [46], MAE3D [19], GL-MAE [20] adopted UNETR [65] as networks architecture. Most other methods, including our MiM, used Swin-UNETR [21] as per the settings of previous works [9]. Table 3 includes the details of the backbone for 3D UNet [37], UNETR [65], and Swin-UNETR [21]. UNETR [65] uses ViT [50] and Swin-UNETR [21] employs the Swin Transformer [45]. For all experiments, we used ViT-Base [50] for UNETR [65] and Swin-Base [45] for Swin-UNETR [21], balancing performance and computational efficiency. These pre-trained encoders were used to initialize the encoder of respective segmentation networks.
Remark. As shown in Table 3, we observed that General SSL methods performed worse than Medical SSL methods. Specifically, SimCLR [29] and MoCov3 [66] achieved only 73.85% and 79.54%, respectively. This is because these methods rely on a large batch size and negative samples to avoid trivial constants, which is impractical for 3D medical images. Additionally, the negative relationship between different images used in SimCLR [29] and MoCov3 [66] is not suitable for 3D medical images. DINO [49] also achieved limited improvements. Our proposed MiM, outperformed MAE-based methods such as MAE3D [19], GL-MAE [20], localMIM [46], HPM [46], and MCMAE [25] by a significant margin. We concluded that General SSL methods are not very suitable for 3D medical images, and it’s crucial to consider the characteristics of 3D medical images when designing SSL methods.
The scratch Swin-UNETR [21] only achieves 80.53% DSC. By pre-training MiM on a 1k unlabeled dataset, we gained a 3.93% improvement with 84.46% DSC, which outperforms existing methods by a clear margin. Among the compared methods, Swin-UNETR [9] and MAE3D [19] achieved the best 82.58% and second-best 82.40% DSC, respectively. Our MiM surpasses these two methods by 1.88% and 2.06% DSC, respectively, which is a clear improvement on this dataset.
It’s worth noting that scaling laws [71] also apply to 3D medical image pre-training. By pre-training with larger scale unlabeled 10k dataset, we observed that Swin-UNETR [9] and MAE3D [19] achieved DSC scores of 83.20% and 83.52%, respectively. Our MiM with 10k achieved a DSC score of 85.41%, which consistently outperformed these two methods significantly. These results suggest that scaling plays an important role in pre-training for 3D medical images and that our MiM method is effective for pre-training on larger datasets.
Qualitative results. MiM was found to improve the completeness of segmentation results, as shown in Fig 6. The results of segmentation using MiM were better than existing methods.
@@ 4.3.2 Comparison on the Unseen datasets
We further conduct experiments on unseen datasets in pre-training, i.e., MM-WHS [60], Spleen [61], Amos22 [63], and Flare22 [62]. The results of these four datasets are shown in Table 4. It can be observed that our MiM consistently outperformed all existing methods with a clear margin, which demonstrates promising generalizability to unseen datasets. Specifically, MiM outperformed existing methods by at least 1.89% DSC on average. By pre-training with a larger scale of unlabeled dataset 10k, Swin-UNETR [9] and MAE3D [19] improved by 0.92% and 0.94% DSC with 86.32% and 86.64%, respectively. Our MiM also gained improvement to 88.34% DSC and outperformed these two methods consistently. MiM also shows label efficiency when finetuning with less labels [9]. Specifically, MiM with 50% label achieved comparable performance with the scratch Swin-UNETR with 100% label with a clear margin.
Table 5: Experiment results on Five CT-based tasks on MSD dataset [61].
Method Pre-training dataset Network Task03 Liver Task06 Lung Task07 Pancreas Task08 Hepatic Vessel Task10 Colon Avg
@@
4.3.3 Comparison on the MSD datasets
To evaluate the generalizability on organ segmentation tasks, we conducted experiments on the five CT-based tasks on the MSD dataset [61], i.e., Task 03 Liver, Task06 Lung, Task07 Pancreas, Task08 Hepatic Vessel and Task10 Colon. Since existing methods didn’t conduct experiments with the same pre-training dataset, we reimplemented these methods for fair comparison. It can be observed in Table 5 that MiM achieves the best average DSC (70.07%) and NSD(78.75%) in all tasks. Since the scratch Swin-UNETR [21] performs better than UNETR [65] in terms of avg DSC (64.84% vs 62.98%) and NSD (73.12% vs 69.08). We further pre-trained UNETR [65] based on MiM for a fair comparison. It can be observed that with MiM pre-training, Swin-UNETR [21] gained an average improvement 5.23% and 5.63% in terms of DSC and NSD, respectively. With UNETR [65] as the network, we observed an average improvement of 10.09% and 9.29% in terms of DSC and NSD, respectively. Moreover, by pre-training with a larger scale of unlabeled dataset 10k, MiM further improved to 70.76% and 79.67% in DSC and NSD, respectively.
4.3.4 Comparison on CC-CCII dataset
To evaluate the generalizability of our MiM on classification task, we fine-tuned it on the CC-CCII [64] dataset and compared its performance with state-of-the-art general and medical SSL methods. Since existing SSL methods didn’t conduct experiments on this dataset, we reproduced the related methods and reported the results. As shown in Table 6, it can be observed that MiM achieved the best performance in terms of ACC and AUC, surpassing all other methods with 93.63% and 99.39%. These findings demonstrate that the learned representation by MiM can be well transferable to classification problems and can be used effectively for medical image classification tasks. With a larger scale of pre-training dataset 10k, our MiM further improved to 94.12% and 99.52% in ACC and AUC, respectively, which shows the scalability of our proposed method when transferring across tasks.
@@ 4.3.4 Comparison on CC-CCII dataset
4.3.5 Comparison on the BraTS 21 dataset
To evaluate the generalizability of our MiM on MRI datasets, we fine-tuned it on the BraST 21 [61] MRI tumor segmentation dataset and compared its performance with state-of-the-art general and medical SSL methods. WT, TC, and ET denote the whole tumor, tumor core, and enhancing tumor, respectively. It can be observed in Table 7 that both SSL methods can improve the performance of the model in segmenting tumors in BrasTS 21 [61] dataset. This is because CT and MRI are often used for the same task but for different purposes, and thus share similar anatomical structures. Therefore, the knowledge learned by SSL methods from unlabeled CT datasets can be transferred to MRI datasets [72, 9]. Our MiM outperformed all other methods by at least 1.34% with 79.28% DSC. By pre-training with a larger unlabeled dataset 10k, our MiM further improved to 79.92%, which shows the scalability of our proposed method when transferring across modalities.
Table 7: Experiment results on BraTS-21 [61].
@@ 4.4 Analysis of our proposed method
All the models were pre-trained on 1k, which were then evaluated on BTCV [53] and MM-WHS [60].
@@ 4.4.1 Ablation study
Loss functions. We conducted comprehensive ablation studies on the BTCV [53] and MM-WHS [60] validation datasets to evaluate the effectiveness of our hierarchical design, focusing on multi-level reconstruction and cross-level alignment components. As shown in Table 8, the multi-level reconstruction losses $L_{R}^{1}$, $L_{R}^{2}$, and $L_{R}^{3}$ substantially enhance model performance by capturing features at multiple granularities from 3D medical images, outperforming traditional single-level mask auto-encoding approaches (Rows 3-4). The importance of proper loss function composition is evident when examining the cross-level alignment loss $L_{C}$ in isolation (Row 2), which resulted in poor convergence with modest performance: 81.92% DSC, 77.11% NSD on BTCV and 89.99% DSC, 74.12% NSD on MM-WHS. This observation indicates that reconstruction losses $L_{R}$ play a fundamental role in establishing robust within-level representations before cross-level alignment $L_{C}$ can be effectively applied. Further analysis reveals that alignment $L_{C}^{2,3}$ between finer levels $x^{2}$ and $x^{3}$ (Row 5) shows modest gains since fine-grained details are primarily addressed by reconstruction losses, whereas alignment $L_{C}^{1,2}$ between coarser levels $x^{1}$ and $x^{2}$ (Row 6) yields more substantial improvements in both DSC and NSD metrics by effectively capturing global context. The optimal performance across both datasets was achieved by combining all reconstruction and cross-level alignment losses (Last row), demonstrating the synergistic relationship between multi-level reconstruction and cross-level alignment in hierarchical representation learning.
Table 8: Evaluation of the loss terms $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$.
@@
${\color[rgb]{0,0,0}\mathcal{L}_{\mathcal{R}}^{1}}$ ${\color[rgb]{0,0,0}\mathcal{L}_{\mathcal{R}}^{2}}$ ${\color[rgb]{0,0,0}\mathcal{L}_{\mathcal{R}}^{3}}$ ${\color[rgb]{0,0,0}\mathcal{L}_{\mathcal{C}}^{12}}$ ${\color[rgb]{0,0,0}\mathcal{L}_{\mathcal{C}}^{23}}$ DSC(%) NSD(%) DSC NSD(%)
@@
Patch Types for Generating subsequent level volumes. During the patchification process (Figure 3), each level-$l$ volume ($x^{l}$) is divided into masked and unmasked patches. We evaluate both patch types for generating the subsequent level-$l+1$ volume ($x^{l+1}$) in our framework. As shown in Table 9, using masked patches from $x^{l}$ to generate $x^{l+1}$ consistently achieves better performance than using unmasked patches from $x^{l}$. This superiority stems from masked patches forcing the model to recover missing information, thereby promoting effective reconstruction of overlapping regions across levels and learning of multi-scale semantic representations. In contrast, unmasked patches directly expose original volume information to the model during the same iteration, effectively creating an information shortcut. This shortcut reduces the learning challenge by allowing the model to simply copy unmasked features rather than learning to infer and reconstruct them, ultimately limiting its ability to capture rich semantic details and develop robust generalization capabilities.
Table 9: Evaluation of the patch types derived from level-l volumes $x^{l}$s utilized for subsequent level volumes $x^{l+1}$s. We report the DSC(%) and NSD(%) on BTCV [53] and MM-WHS [60].
Patch Types for Generating $x^{l+1}$ BTCV MM-WHS
@@
Unmasked patches from $x^{l}$ 82.18 76.22 89.33 73.65
Masked patches from $x^{l}$ 84.66 82.11 91.04 76.65
Negative pairs for $\mathcal{L}_{\mathcal{C}}$. In Eq. 6, we employ infoNCE loss[29] as the default choice. This loss maximizes the similarity between the cross-level images and pushes away the negative samples. Another loss function is BYOL cosine loss [73]. The primary distinction is whether to use negative samples. As in Table 10, negative samples aid in learning better representations [29], and thus, we include them in our default choices.
Table 10: Evaluation of $\mathcal{L}_{C}$ whether using negative samples.
@@
Efficiency of MiM. This study adopts a hybrid convolution-transformer backbone that integrates convolutional blocks with transformer layers. By incorporating convolutional blocks, the architecture enhances inductive bias learning and enables the reuse of multi-scale features, effectively supporting hybrid representation learning [25]. This design enhances the transformer’s capability to process 3D medical images with greater efficiency and effectiveness. As such, this backbone is a foundational component of our proposed framework. Table 11 presents a comparative analysis of computation costs i.e. flops and times during the pre-training period across various methodologies. Our evaluation highlights the computational efficiency of MAE-based approaches when coupled with UNETR, as they leverage only the unmasked tokens from masked 3D medical images. In contrast, alternative MAE-based methods naively cooperating with hybrid backbones such as Swin-UNETR, necessitate processing all tokens. Our proposed method extends the hybrid backbone architecture [25] to 3D medical images in the pre-training stage, achieving significantly superior performance while maintaining computational efficiency.
Table 11: Evaluation of model complexity during pre-training and subsequent performance on downstream tasks.
Method Networks Tokens flops(G) Pre-training duration(h) DSC(%)
All Partial BTCV MM-WHS
@@ 4.4.2 Hyper-parameter analysis
Weights of $\mathcal{L}_{\mathcal{R}}$ and $\mathcal{L}_{\mathcal{C}}$. In Eq. 10, the loss function of MiM consists of two parts. Thus, in Fig 8, we increased the value of $\alpha$ The optimal value of $\alpha$ was $1e^{-1}$ by observation. Further scaling up the weight of cross-level alignment did not bring any additional benefits. This may be because the magnitude of the cross-level alignment function is much larger than the magnitude of the reconstruction loss, resulting the ignorance of reconstruction process.
Figure 7: Evaluation of the loss weight $\alpha$ for $\mathcal{L_{C}}$.
Figure 8: Evaluation of the number of level $L$.
[…]
Table 12: Evaluation of model with different learning process.
Learning process Weight of Level-l Datasets
$\mathcal{L}_{R}^{1}$ $\mathcal{L}_{R}^{2}$ $\mathcal{L}_{R}^{3}$ BTCV MM-WHS
@@
Numbers of multi-levels $L$. Our adoption of a three-level architecture is supported by both empirical evidence and practical considerations. Empirically, as demonstrated in Fig 9, our multi-level strategy preserves anatomical structures through carefully selected resolutions: Level-1 (384$\times$384$\times$192) captures broad contextual information, Level-2 (96$\times$96$\times$96) extracts intermediate features, and Level-3 (16$\times$16$\times$16) retains fine details. Through extensive experimentation shown in Fig 8, we investigated the impact of varying the number of levels $L$ and found that $L=3$ yields optimal performance. Adding a fourth level would result in an extremely low resolution ($1\times 1\times 1$), which, while potentially simplifying reconstruction, leads to suboptimal performance due to excessive information loss, consistent with findings in similar architectures[17, 39].
Figure 9: Illustration of our multi-level volume generation strategy. The framework processes volumes at three hierarchical levels with decreasing patch sizes: Level-1 (96 px patches), Level-2 (16 px patches), and Level-3 (3x3 patches). This design captures anatomical information at complementary scales while maintaining computational efficiency. Further subdivision beyond Level-3 would result in patches too small to preserve meaningful anatomical features. Gray patches indicate masked regions, while color intensity (light to dark) represents the progression from coarse to fine granularity across three hierarchical levels.